# Quantitative Understanding of PDF Fits and their Uncertainties

### v1

@@ 3.2.1 NTK at initialisation
It is argued in the literature that, in the large-width limit, the variance of the NTK over the set of replicas tends to zero with the width of the hidden layers (see, e.g., [Roberts:2021fes, Chiefa:2026TBA]). In order to quantify the variation of the NTK, we start by computing the Frobenius norm of the NTK over an ensemble of networks for different architectures. For each architecture, we consider the standard deviation of the norm as a statistical estimator of the variations of the NTK. The result is displayed in Fig. 6. Even though the Frobenius norm is a coarse indicator of the variations of the NTK, the figure shows clearly that the variance of the norm becomes smaller with the size of the network, which is consistent with the theoretical expectation that the NTK should not fluctuate for infinite-width networks [^7] Note that, in addition to the scaling $\mathcal{O}(1/n)$ theoretically predicted for large networks, the uncertainty bands include bootstrap errors due to the finite size of the ensemble. Using an ensemble of 100 replicas, the bootstrap error on the standard deviation is $\sim 10\%$..
@@
A more quantitative description of the NTK at initialisation is provided by its spectrum, which is shown in Fig. 7 for four different architectures. Inspecting the figure, we see that the spectrum of the NTK is heavily hierarchical, and only few eigenvalues are actually non-zero [^8] Note that, due to the large difference in magnitude of the eigenvalues, the finite precision used in our codes introduces noise in the decomposition, so that small eigenvalues should be effectively considered zero. We discuss the cut-off tolerance in Appendix C.. Such a hierarchy in the eigenvalues means that only a small subset of active directions can inform the network during training, as it will be discussed later. Note that, at least at initialisation, these observations do not depend on the architecture: the eigenvalues in Fig. 7 are mostly independent of the size of the network. Even though the logarithmic scale on the vertical axis may hide some small variations, it is clear that most eigenvalues remain constant within the error bars. On the other hand, the logarithmic scale emphasises that there are several orders of magnitude between eigenvalues for a given architecture; that hierarchical structure does not depend on the architecture. There is a downward fluctuation of the third eigenvalue for the largest architecture that we considered, but we do not have any evidence that this drop is a physical feature of the system, rather than a fluctuation. Finally, the variance of the set of eigenvalues over replicas decreases with increasing size, as expected.
@@ 3.2.2 NTK During Training
We now discuss the behaviour of the NTK during training. To this end, we are going to adopt the so-called closure tests developed by the NNPDF collaboration. A closure test uses synthetic data, generated using a known set of PDFs, to train the neural network. The PDFs used for generating the data are called here input PDFs. The results of the training are then compared to the known input PDFs; the performance of the training algorithm and the NN architecture are assessed by quantifying the comparison between trained PDFs and input PDFs. Following the original presentation in Ref. [NNPDF:2014otw], we distinguish three levels of closure tests, which are defined by the complexity of the data used to train the NNs. We use the standard NNPDF nomenclature and refer to these three levels as level-0 (L0), level-1 (L1), and level-2 (L2) closure tests, and we denote the input PDFs used to generate the data as $f^{\rm in}$. The definitions of these three levels of data are given in Appendix A.
@@
for the three different datasets, L0, L1, and L2. Inspecting the plot reveals that the NTK undergoes significant changes during the initial phase of training, with the relative variation $\delta\Theta_{t}$ reaching values as high as $6\%$. This indicates that our settings differ from the standard picture of lazy training in the context of very wide networks, as discussed, e.g., in Refs. [jacot2018neural, Roberts:2021fes, lee2019wide], where the NTK is expected to be independent of the flow time $t$. Remarkably, we do not observe a dependence on how data have been generated, indicating that the NTK dynamics is basically unaffected by the noise that affects the data.
After this initial phase – corresponding approximately to the first 20,000 epochs in our experiment – the NTK tends to stabilize. These two regions will be referred to as the rich and lazy training regimes, respectively, in keeping with the standard terminology adopted in the literature (see, e.g., Ref. [fort2020dlvk] where two similar regimes were also identified). We do not comment any further on the implications of the lazy regime, and postpone the discussion to Sec. 4.1.1.
@@ 3.2.3 Eigenvectors and Alignment of the NTK
The blue rectangle in the top right corner of the matrix shows that the eigenvectors of the NTK corresponding to the largest eigenvalues are orthogonal to the eigenvectors of $M$ that are in the kernel of $M$, i.e., the directions that do not contribute to the observables. It is useful to remember that the largest eigenvalues of the NTK correspond to the directions that are orthogonal to $\ker\Theta$, i.e., the directions that are learnable during the training process. In order to have a robust training process, we expect these learnable directions to align with the directions that actually contribute to the loss functions, i.e., the ones corresponding to the largest eigenvalues of $M$. Consistently with this intuition, we see that the size of this blue rectangle increases with training time. In particular, it is clear from our plot that it becomes deeper by the onset of the lazy training regime: more of the learnable directions – the features that the network can learn – are aligned with the directions that contribute most to the observables.
A similar analysis can also be performed by studying the alignment of the eigenvectors of the NTK with the input function used to generate the data. In Fig. 13, we show the cosine of the angle between the eigenvectors $z^{(k)}$ and the input $f^{\mathrm{in}}$ for two subsets of eigenvectors, in the right and left panel respectively. In these two panels, different patterns can be observed. Firstly, the first four angles vary considerably during training, consistent with the observations made in the figures before. Furthermore, we see that other than changing the dominant eigenvectors, the NTK is also activating the other directions that were subdominant in the first stage of training, in agreement with the observations coming from Figs. 10-12. By the onset of the lazy training regime, the input function lies entirely in the subspace spanned by the first four eigenvectors of the NTK. This confirms that the NTK has aligned its features with the physically relevant directions of the problem.
@@
Figure 13: Alignment of the eigenvectors of the NTK with the input function $f^{\rm in}$ used to generate the L2 data, measured in terms of $\cos\theta^{(k)}=(z^{(k)},f^{\rm in})/\|f^{\rm in}\|$. In the left panel, the first four eigenvectors of the NTK are shown, while the right panel shows the remaining eigenvectors up to $k=7$. Note that $f^{\mathrm{in}}$ lies entirely in the subspace spanned by the first four eigenvectors of the NTK by the onset of the lazy training regime. We see that the NTK has aligned its features with the physically relevant directions of the problem.
A complementary picture is displayed in Fig. 14. Here, we show the eigenvectors of the product of matrices $\Theta_{t}M$, labelled with $q^{(i)}$, at different training times and as functions of the $x$-grid. The product of these two matrices will become relevant in Sec. 4, where the solution of the lazy training equations is expressed as a linear combination of these eigenvectors. Together with the eigenvectors, we also show the output of the trained neural network at the corresponding training time. From these plots, we see that as the training progresses, the shape of the eigenvectors becomes more structured in order to reproduce the output function. Again, this conclusion supports the observations made previously in various occasions, that during training the neural network is changing its internal representation and the NTK encodes this information.
@@
Figure 14: First five eigenvectors of the combined matrix $H=\Theta M$, as in Eq. (34), at different training times and as function of the input $x$-grid. We also show the output of the network at the same training time, which is displayed in gray. L1 data is used.
@@ 4 Lazy Training in NNPDF
In the previous section we presented an empirical study of the training dynamics through the lens of the NTK. We observed that the NTK is able to capture the main features of the training process, and that its time evolution is characterised by a rapid initial transient, followed by a slower evolution during the rest of the training. We now turn our attention on this last stage of the training, where the NTK has stabilised and becomes approximately constant. In doing so, we build upon the results presented in Refs. [jacot2018neural, lee2019wide] and extend them to the case of NNPDF. In the following, we derive the analytical solution of the flow equation, which allows us to write an explicit expression for the trained field as a function of the field at initialisation and the data.
@@ 4.1.1 Solution of the Flow Equation
In Fig. 15 we show the evolution during training of the first five eigenvalues of $H^{\perp}$ for the three different closure datasets. It should not come as too much of a surprise that the eigenvalues of $H^{\perp}$ have a similar behaviour to those of the NTK (see Fig. 9), from which they are constructed. However, we see that $h^{(i)}$ are larger by around three orders of magnitude than the NTK eigenvalues.
@@
The solution in Eq. (56) is the main result of this section. It shows that the training process can be described as the sum of a linear transformation of the initial fields $f_{0,\alpha}$, and a linear transformation of the data $Y_{I}$. The two transformations depend on the flow time $t$ and are given by the evolution operators $U(t)$ and $V(t)$. Fig. 16 compares the analytical solution with the trained function at the end of training, for different choices of the frozen NTK. The NN is trained using the numerical GD until $t_{\rm ref}$, at which point the NTK is frozen. The evolution time $t$ used in the analytical solution is the difference between the total training time and $t_{\rm ref}$; the intial condition for the analytical solution is the trained solution at $t_{\rm ref}$. Central value and uncertainty bands are obtained by computing the analytical solution for each replica of the initial condition and frozen NTK.  [^10] If not stated otherwise, in this section central values and uncertainties are always computed as ensemble averages across replicas. The figures agree with the expectation; the closer $t_{\rm ref}$ is to the onset of the lazy regime, the better the agreement between the analytical solution and the trained function.
@@ 4.2.4 Error decomposition
The analytical expression for the covariance matrix, Eq. (76), allows us to monitor the relative size of the three contributions as training proceeds. For a properly trained ensemble of networks, the covariance of the trained fields should be dominated by the statistical error on the data. We show the diagonal entries of the decomposition of the error budget in Fig. 22, for different frozen NTKs (rows) and different training epochs (columns), using L2 data as before. In general, we observe that towards the end of the training process the contribution from the data $C_{t}^{(YY)}$ (orange band) becomes dominant with respect to the contribution coming from the initial condition $C_{t}^{(00)}$ (blue band). We also see that the suppression of the initial condition is more severe and happens earlier when the frozen NTK is taken at later stages of training.
In order to study the dependence of the error decomposition on the initial condition, we repeat the same analysis for the case of scaled input $f(x,\log x)$, as described in Sec. 2. Hence, the frozen NTK is taken from an ensemble of trained networks with input layer as in Eq. (8). The initial condition $f_{0}$ of the analytical solution is drawn from the same prior distribution as the trained solution, and corresponds to the orange curve in Fig. 5. The resulting error decomposition is shown in Fig. 23 for L2 data, where panels are ordered as in Fig. 22. Inspecting the figures, we observe that now the contribution of the initial condition becomes dominant in the small-$x$ region, even for large training times and irrespective of the epoch at which the NTK is frozen. This result reflects the behaviour of the prior distribution at small-$x$, where indeed error bands are significantly enlarged with respect to the case of linear input. Interestingly, even the contribution from the data, $C^{(YY)}_{t}$, increases at small-$x$ towards larger training times. This can be explained by observing that, despite not being explicitly dependent on the initial condition $f_{0}$, the evolution operator $V(t)$ is constructed from a frozen NTK that has encoded the dependence on the architecture through the training process (see Eq. (62)). That the difference between Figs. 22 and 23 is primarily localised at small-$x$ showcases that, for the region left uncovered by the data, the methodology is not able to suppress the dependence on the initial condition. In fact, where the information from the data is available (corresponding roughly to $x\gtrsim 0.01$), the dependence on the initial condition lessens as the analytical solution evolves (left to right). This reduction occurs more rapidly for frozen NTKs taken at later stages of training (top to bottom), showing that there is a non-trivial interplay between the information provided by the data and that acquired by the NTK.
@@ 4.3 Bias-Variance Decomposition
Finally, we can use the analytical solution to compute the bias and variance of the trained ensemble of networks. Following the definition of Ref. [NNPDF:2021njg], we define the bias as
@@
Continuing in this vein and following Ref. [NNPDF:2021njg], we define the variance as
@@
Bias and variance, together with their decompositions introduced above, are shown in Figs. 24 and 25, respectively. We show these curves as functions of the training time for different frozen NTKs and for L0, L1 and L2 data. We now discuss these figures in turn.
@@
Figure 26: Comparison of the total contribution of bias and variance in the space of the data as a function of the training time for L0 (top), L1 (middle) and L2 (bottom) data. The curves are the same as in Figs. 24 and 25, displayed here together. Each column corresponds to a different frozen NTK. In first column, the bias is off-scale. As in Fig. 25, the vertical dashed lines in grey indicate the inverse of the eigenvalues of $H^{\perp}$, $1/h^{(i)}$.

### v2

@@ 3.2.1 NTK at initialisation
It is argued in the literature that, in the large-width limit, the variance of the NTK over the set of replicas tends to zero with the width of the hidden layers (see, e.g., [33, 16]). In order to quantify the variation of the NTK, we start by computing the Frobenius norm of the NTK over an ensemble of networks for different architectures. For each architecture, we consider the standard deviation of the norm as a statistical estimator of the variations of the NTK. The result is displayed in Fig. 6. Even though the Frobenius norm is a coarse indicator of the variations of the NTK, the figure shows clearly that the variance of the norm becomes smaller with the size of the network, which is consistent with the theoretical expectation that the NTK should not fluctuate for infinite-width networks. [^7] Note that, in addition to the scaling $\mathcal{O}(1/n)$ theoretically predicted for large networks, the uncertainty bands include bootstrap errors due to the finite size of the ensemble. Using an ensemble of 100 replicas, the bootstrap error on the standard deviation is $\sim 10\%$.
@@
A more quantitative description of the NTK at initialisation is provided by its spectrum, which is shown in Fig. 7 for four different architectures. Inspecting the figure, we see that the spectrum of the NTK is heavily hierarchical, and only few eigenvalues are actually non-zero. [^8] Note that, due to the large difference in magnitude of the eigenvalues, the finite precision used in our codes introduces noise in the decomposition, so that small eigenvalues should be effectively considered zero. We discuss the cut-off tolerance in Appendix C. Such a hierarchy in the eigenvalues means that only a small subset of active directions can inform the network during training, as it will be discussed later. Note that, at least at initialisation, these observations do not depend on the architecture: the eigenvalues in Fig. 7 are mostly independent of the size of the network. Even though the logarithmic scale on the vertical axis may hide some small variations, it is clear that most eigenvalues remain constant within the error bars. On the other hand, the logarithmic scale emphasises that there are several orders of magnitude between eigenvalues for a given architecture; that hierarchical structure does not depend on the architecture. There is a downward fluctuation of the third eigenvalue for the largest architecture that we considered, but we do not have any evidence that this drop is a physical feature of the system, rather than a fluctuation. Finally, the variance of the set of eigenvalues over replicas decreases with increasing size, as expected.
@@ 3.2.2 NTK During Training
We now discuss the behaviour of the NTK during training. To this end, we are going to adopt the so-called closure tests developed by the NNPDF collaboration. A closure test uses synthetic data, generated using a known set of PDFs, to train the neural network. The PDFs used for generating the data are called here input PDFs. The results of the training are then compared to the known input PDFs; the performance of the training algorithm and the NN architecture are assessed by quantifying the comparison between trained PDFs and input PDFs. Following the original presentation in Ref. [8], we distinguish three levels of closure tests, which are defined by the complexity of the data used to train the NNs. We use the standard NNPDF nomenclature and refer to these three levels as level-0 (L0), level-1 (L1), and level-2 (L2) closure tests, and we denote the input PDFs used to generate the data as $f^{\rm in}$. The definitions of these three levels of data are given in Appendix A.
@@
for the three different datasets, L0, L1, and L2. Inspecting the plot reveals that the NTK undergoes significant changes during the initial phase of training, with the relative variation $\delta\Theta_{t}$ reaching values as high as $6\%$. This indicates that our settings differ from the standard picture of lazy training in the context of very wide networks, as discussed, e.g., in Refs. [28, 33, 30], where the NTK is expected to be independent of the flow time $t$. Remarkably, we do not observe a dependence on how data have been generated, indicating that the NTK dynamics is basically unaffected by the noise that affects the data.
After this initial phase – corresponding approximately to the first 20,000 epochs in our experiment – the NTK tends to stabilize. These two regions will be referred to as the rich and lazy training regimes, respectively, in keeping with the standard terminology adopted in the literature (see, e.g., Ref. [23] where two similar regimes were also identified). We do not comment any further on the implications of the lazy regime, and postpone the discussion to Sec. 4.1.1.
@@ 3.2.3 Eigenvectors and Alignment of the NTK
The blue rectangle in the top right corner of the matrix shows that the eigenvectors of the NTK corresponding to the largest eigenvalues are orthogonal to the eigenvectors of $M$ that are in the kernel of $M$, i.e., the directions that do not contribute to the observables. It is useful to remember that the largest eigenvalues of the NTK correspond to the directions that are orthogonal to $\ker\Theta$, i.e., the directions that are learnable during the training process. In order to have a robust training process, we expect these learnable directions to align with the directions that actually contribute to the loss functions, which are the ones corresponding to the largest eigenvalues of $M$. Consistently with this intuition, we see that the size of this blue rectangle increases with training time. In particular, it is clear from our plot that it becomes deeper by the onset of the lazy training regime: more of the learnable directions – the features that the network can learn – are aligned with the directions that contribute most to the observables.
@@
Figure 13: Reconstruction error $C_{t}(r)$ as defined in Eq. (38) as a function of the training time and for different numbers of eigenvectors $r$. Note that $f^{\rm in}$ lies entirely in the subspace spanned by the first four eigenvectors of the NTK by the onset of the lazy training regime. We see that the NTK has aligned its features with the physically relevant directions of the problem.
We remark that the eigenvectors of the NTK form an orthonormal basis in $\mathbb{R}^{N_{\mathrm{grid}}}$ at any epoch of the training process. We have also shown that only a subset of these eigenvectors contribute to the training dynamics. We may thus wonder how expressive the vectors belonging to the subspace orthogonal to the kernel are, i.e., how well the internal representation of the neural network can reconstruct the input function $f^{\rm in}$ used to generate the data. In order to quantify this aspect, we define a new figure of merit
$C_{t}(r)=1-\sum_{k=1}^{r}\frac{(z^{(k)},f^{\rm in})^{2}}{\|f^{\rm in}\|^{2}}\,$, (38)
which measures the reconstruction error of the input function $f^{\rm in}$ when projected onto the subspace spanned by the first $r$ eigenvectors of the NTK at training time $t$. We show $C_{t}(r)$ as a function of the training time and for different choices of $r$ in Fig. 13. Inspecting the figure, we see that in the early stages of training the reconstruction error does not change significantly with time. This behaviour is shared across all values of $r$. In fact, in this first phase the NTK has not yet found a suitable internal representation and the inclusion of more eigenvectors corresponding to yet undiscovered directions – those associated with a small eigenvalue as in Fig. 9 – does not result in an improvement of the reconstruction of $f^{\rm in}$. Conversely, once the onset of lazy training is approached, we identify two distinct behaviours depending on the number of used eigenvectors. If we include the eigenvectors up to the modes discovered during training (e.g., $r=6$ and $r=7$ in the figure), the reconstruction error drops significantly and remains almost constant throughout the rest of the learning process. On the other hand, if we consider fewer eigenvectors (e.g., $r=4$ and $r=5$ in the figure), the reconstruction error becomes larger and grows indefinitely with training time. This results show us that the subset of eigenvectors orthogonal to $\ker\Theta$ is capable of providing an effective lower dimensional basis, provided that the new modes discovered by the NTK during training are included. Note that the reconstruction of the input function improves as long as the included eigenvectors do not belong to $\ker\Theta$ – being noise, they would not bring any physical information.
@@
Figure 14: First five eigenvectors of the NTK at different training times and as function of the input $x$-grid. We also show the output of the network at the same training time, which is displayed in gray. L1 data is used.
A complementary picture is displayed in Fig. 14. Here, we show the eigenvectors of the NTK at different training times as functions of the $x$-grid, denoted by $z^{(i)}$. Together with the eigenvectors, we also show the output of the trained neural network at the corresponding training time. From these plots, we see that as the training progresses, the shape of the eigenvectors becomes more structured in order to reproduce the output function. Again, this conclusion supports the observations made previously on various occasions, that during training the neural network is changing its internal representation and the NTK encodes this information.
@@ 4 Lazy Training in NNPDF
In the previous section we presented an empirical study of the training dynamics through the lens of the NTK. We observed that the NTK is able to capture the main features of the training process, and that its time evolution is characterised by a rapid initial transient, followed by a slower evolution during the rest of the training. We now turn our attention on this last stage of the training, where the NTK has stabilised and becomes approximately constant. In doing so, we build upon the results presented in Refs. [28, 30] and extend them to the case of NNPDF. In the following, we derive the analytical solution of the flow equation, which allows us to write an explicit expression for the trained field as a function of the field at initialisation and the data.
@@ 4.1.1 Solution of the Flow Equation
Figure 16: First five eigenvectors of the combined matrix $H=\Theta M$, as in Eq. (34), at different training times and as a function of the input $x$-grid. We also show the output of the network at the same training time, which is displayed in gray. L1 data is used.
In Fig. 15 we show the evolution during training of the first five eigenvalues of $H^{\perp}$ for the three different closure datasets. In Fig. 16 we show the first five eigenvectors of $H$, denoted as $q^{(i)}$, at different training times as functions of the $x$-grid. It should not come as too much of a surprise that the eigenvalues and eigenvectors of $H^{\perp}$ have a similar behaviour to those of the NTK (see Figs. 9 and 14), from which they are constructed. However, we see that $h^{(i)}$ are larger by around three orders of magnitude than the NTK eigenvalues.
@@
The solution in Eq. (57) is the main result of this section. It shows that the training process can be described as the sum of a linear transformation of the initial fields $f_{0,\alpha}$, and a linear transformation of the data $Y_{I}$. The two transformations depend on the flow time $t$ and are given by the evolution operators $U(t)$ and $V(t)$. Fig. 17 compares the analytical solution with the trained function at the end of training, for different choices of the frozen NTK. The NN is trained using the numerical GD until $t_{\rm ref}$, at which point the NTK is frozen. The evolution time $t$ used in the analytical solution is the difference between the total training time and $t_{\rm ref}$; the intial condition for the analytical solution is the trained solution at $t_{\rm ref}$. Central value and uncertainty bands are obtained by computing the analytical solution for each replica of the initial condition and frozen NTK. [^10] Unless stated otherwise, in this section central values and uncertainties are always computed as ensemble averages across replicas. The figures agree with the expectation; the closer $t_{\rm ref}$ is to the onset of the lazy regime, the better the agreement between the analytical solution and the trained function.
@@ 4.2.4 Error decomposition
The analytical expression for the covariance matrix, Eq. (77), allows us to monitor the relative size of the three contributions as training proceeds. For a properly trained ensemble of networks, the covariance of the trained fields should be dominated by the statistical error on the data. We show the diagonal entries of the two contributions $C_{t}^{(00)}$ (blue band) and $C_{t}^{(YY)}$ (orange band) to the error budget in Fig. 23, for different frozen NTKs (rows) and different training epochs (columns), using L2 data as before. We do not show the mixed term $C_{t}^{(0Y)}$ since it is negligible with respect to the other two contributions – the two sources of uncertainty are largely uncorrelated. In general, we observe that towards the end of the training process the contribution from the data $C_{t}^{(YY)}$ becomes dominant with respect to the contribution coming from the initial condition $C_{t}^{(00)}$. We also see that the suppression of the initial condition is more severe and happens earlier when the frozen NTK is taken at later stages of training.
In order to study the dependence of the error decomposition on the initial condition, we repeat the same analysis for the case of scaled input $f(x,\log x)$, as described in Sec. 2. Hence, the frozen NTK is taken from an ensemble of trained networks with input layer as in Eq. (8). The initial condition $f_{0}$ of the analytical solution is drawn from the same prior distribution as the trained solution, and corresponds to the orange curve in Fig. 5. The resulting error decomposition is shown in Fig. 24 for L2 data, where panels are ordered as in Fig. 23. Inspecting the figures, we observe that now the contribution of the initial condition becomes dominant in the small-$x$ region, even for large training times and irrespective of the epoch at which the NTK is frozen. This result reflects the behaviour of the prior distribution at small-$x$, where indeed error bands are significantly enlarged with respect to the case of linear input. Interestingly, even the contribution from the data, $C^{(YY)}_{t}$, increases at small-$x$ towards larger training times. This can be explained by observing that, despite not being explicitly dependent on the initial condition $f_{0}$, the evolution operator $V(t)$ is constructed from a frozen NTK that has encoded the dependence on the architecture through the training process (see Eq. (63)). That the difference between Figs. 23 and 24 is primarily localised at small-$x$ showcases that, for the region left uncovered by the data, the methodology is not able to suppress the dependence on the initial condition. In fact, where the information from the data is available (corresponding roughly to $x\gtrsim 0.01$), the dependence on the initial condition lessens as the analytical solution evolves (left to right). This reduction occurs more rapidly for frozen NTKs taken at later stages of training (top to bottom), showing that there is a non-trivial interplay between the information provided by the data and that acquired by the NTK.
@@ 4.3 Bias-Variance Decomposition
Finally, we can use the analytical solution to compute the bias and variance of the trained ensemble of networks. Following the definition of Ref. [9], we define the bias as
@@
Continuing in this vein and following Ref. [9], we define the variance as
@@
Bias and variance, together with their decompositions introduced above, are shown in Figs. 25 and 26, respectively. We show these curves as functions of the training time for different frozen NTKs and for L0, L1 and L2 data. We now discuss these figures in turn.
@@
Figure 27: Comparison of the total contribution of bias and variance in the space of the data as a function of the training time for L0 (top), L1 (middle) and L2 (bottom) data. The curves are the same as in Figs. 25 and 26, displayed here together. Each column corresponds to a different frozen NTK. In first column, the bias is off-scale. As in Fig. 26, the vertical dashed lines in grey indicate the inverse of the eigenvalues of $H^{\perp}$, $1/h^{(i)}$.